# Quantifying resilience for distribution system customers with SALEDI

### v1

@@ IV-A Why the logarithm in SALEDI is needed
is a Seemingly Plausible Average Large Event Duration Index that is the mean number of normalized CMIp for large events. SPALED is ALED without the logarithm. We now quantify the variability of SPLEDI and SPALED in order to establish that practical estimates of SPLEDI and SPALED from data require too many samples.
@@
Let $Y$ be the random variable for the normalized event CMIp $M/M_{\rm large}$. We will take $Y\sim P_{b}$ or $Y\sim LN_{b}$. According to (24) and (22), SPALED is the mean of $N_{\rm large}$ samples of $Y$, where $N_{\rm large}$ has mean $\overline{n}_{\rm large}$, and SPLEDI is SPALED multiplied by annual frequency $f_{\rm large}$. Then, using Appendix B,
${\rm RSE}[{\rm SPLEDI}]$ $=\frac{\sqrt{1+{\rm RSE}[Y]^{2}}}{\sqrt{\overline{n}_{\rm large}}}$ (27)
${\rm RSE}[{\rm SPALED}]$ $=\frac{{\rm RSE}[Y]}{\sqrt{\overline{n}_{\rm large}}}$ (28)
We can compare the variability of SPLEDI (27) with the variability of SALEDI (12) for the same number of samples $\overline{n}_{\rm large}$ by comparing $\sqrt{1+({\rm RSE}_{{\rm Pb}})^{2}}$ or $\sqrt{1+({\rm RSE}_{{\rm LNb}})^{2}}$ with $\sqrt{2}=1.4$. That is, we compare the numerators of (27) and (12). Table I shows ${\rm RSE}_{{\rm Pb}}$ and ${\rm RSE}_{{\rm LNb}}$ for each utility.
For approximating the bounded distribution of CMIp with $Pb$, Table I shows that utilities 1,2,3,4 have $\sqrt{1+({\rm RSE}_{{\rm Pb}})^{2}}$ an order of magnitude greater than $\sqrt{2}$. Then for the same accuracy of an estimate of the mean, SPLEDI requires at least two orders of magnitude more large events than SALEDI, making it impractical as a resilience metric calculated from observed data. For example, if SALEDI requires 2 years of data to accumulate enough large events for reasonable statistical accuracy, then SPLEDI requires $\sim$200 years. ${\rm RSE}_{{\rm Pb}}=13$ for utility 5 which is almost an order of magnitude greater than $\sqrt{2}=1.4$ and requires more data by a factor of $(\sqrt{1+({\rm RSE}_{{\rm Pb}})^{2}}/\sqrt{2})^{2}=(1+13^{2})/2=85$.
For approximating the bounded distribution of CMIp with $LNb$, Table I shows that utilities 1,2,3 have $\sqrt{1+({\rm RSE}_{{\rm LNb}})^{2}}$ at least an order of magnitude greater than $\sqrt{2}$, so that SPLEDI requires at least two orders of magnitude more large events than SALEDI for the same accuracy of estimates of the mean. Utility 4 requires more data by a factor of $(1+9^{2})/2=41$, and utility 5 requires more data by a factor of $(1+2.7^{2})/2=4$ for the same number of large events, $\overline{n}_{\rm large}$.
There is a similar conclusion comparing the variability of SPALED (28) with the variability of ALED (13), since ${\rm RSE}_{{\rm Pb}}$ and ${\rm RSE}_{{\rm LNb}}$ are an order of magnitude greater than 1 except that ${\rm RSE}_{{\rm LNb}}=8.7$ and ${\rm RSE}_{{\rm LNb}}=2.7$ for utilities 4 and 5. Even if SPALED could in theory be computed from very large quantities of data, note that SPALED, since it is a mean, cannot be interpreted as a typical value due to the heavy tails.
@@
IV-B Comparing the use of SALEDI and SAIDI
SAIDI without major event days and SALEDI quantify different aspects of CMIp. SAIDI without major event days focuses on the CMIp of common individual outages and quantifies reliability by excluding major event days, whereas SALEDI focuses on the CMIp of large events and quantifies resilience.
SAIDI with major event days measures the total impact of all sustained outages experienced by customers in a year. Although SAIDI with major event days combines in some way reliability and resilience, it does not give a clear indication of either reliability or resilience because it varies erratically with the large variations in the CMIp of the large events. Moreover, it is necessary and desirable to separately monitor reliability and resilience with different metrics. While improvements to reliability can quite often also improve resilience, they are distinct because the scale and urgency and risk of large event outages and restoration differ so much. Therefore proposed upgrades should be evaluated separately for reliability and resilience.
@@ IV-C CVAR interpretations
Therefore ALED, which works well as an index, is a logarithmic form of CVAR [^11] However, ${\rm CVAR}_{\rm lnM}$ differs from ALED in the definition of the threshold $M_{\rm large}$ for large events. For ${\rm CVAR}_{\rm lnM}$, $M_{\rm large}$ is VAR, the value at a given quantile of normalized event CMIp, whereas for ALED, $M_{\rm large}$ is chosen at the lowest point of the linear region of the exceedance function.. Moreover, (5) shows that
@@ IV-D Heavy tails and extrapolating the trend of extreme events
The bounded Pareto distribution (25) linearly extrapolates the slope of the exceedance curve tail beyond the maximum observed normalized CMIp $M_{\rm maxobs}/M_{\rm large}$ up to the maximum normalized CMIp $M_{\rm max}/M_{\rm large}$ for the largest possible blackout. This allows the likely performance of the metrics to be evaluated while including the unobserved future blackouts of the highest customer impact. The linear extrapolation of the trend is somewhat uncertain, but the extrapolation nevertheless seems more reasonable than ignoring the even more extreme events indicated by the observed trend—it seems unlikely that the largest blackout has already been observed in the data. It is the case that sums of samples of heavy-tailed distributions are dominated by the maximum sample that occurs [^12] This is called a “catastrophe principle” and proved in [16, chapter 3].. For example, both SAIDI with major event days and SPLEDI add up normalized CMIp, and suffer high variability due to this effect. High-impact blackouts with extremely high CMIp values are rare but inevitable due to heavy tails. This emphasizes the importance, despite its inherent uncertainties, of trying to estimate the extrapolated trend governing the most extreme events to encompass these extreme events in the resilience index. The need to obtain this linear extrapolation is one reason that we confine the large events to the linear region of the tail. We assume the existence of a suitable linear region across a sufficient range of the largest observed blackouts, because it is present in the data we analyze here.
@@ V Practical details of SALEDI
V-A Average number of large events $\overline{n}_{\rm large}$
Suppose that $n_{\rm large}^{\rm all}$ large events are observed in all the outage data over $n_{\rm year}^{\rm all}$ years. We estimate the annual frequency of large events using all the observed years of data as
$f_{\rm large}^{\rm all}=n_{\rm large}^{\rm all}/n_{\rm year}^{\rm all}$ (36)
$\begin{aligned} \text{Then }&\hskip 56.9055pt\overline{n}_{\rm large}=f_{\rm large}^{\rm all}n_{\rm year}\end{aligned}$ (37)
is the average number of large events in $n_{\rm year}$ years.
V-B Choosing the large event threshold $M_{\rm large}$
SALEDI is defined over events whose CMIp $M_{i}$ exceeds the $M_{\rm large}$ threshold. We want to select an $M_{\rm large}$ that satisfies the following criteria:
  • $M_{\rm large}$ should be small enough to give sufficient large events to keep the variability ${\rm RSE}_{\rm SAL}$ small.
@@ V-B Choosing the large event threshold $M_{\rm large}$
    Select all values above $M_{0}$, i.e., $S=\{M_{i}:M_{i}\geq M_{0}\}$
@@
    $\alpha_{0}=(1/|S|)\sum_{i=1}^{|S|}{\rm ln}\{M_{i}/M_{0}:M_{i}\geq M_{0}\}$
    where $|S|$ is the number of values in the selected data.
@@
    Calculate Kolmogorov-Smirnov distance $D$ between the empirical CDF of the selected data $F_{S}(M)$ and the CDF $F_{P_{0}}(M)$ of a Pareto distribution with parameter $\alpha_{0}$:
    $D=\underset{M_{i}\geq M_{0}}{\rm max}|F_{S}(M_{i})-F_{P_{0}}(M_{i})|$
@@
V-C Accuracy and number of years $n_{\rm year}$
$M_{\rm large}$ is calculated once and then stays fixed for the rest of the calculations. Given a fixed $M_{\rm large}$, the statistical variability of SALEDI is governed by the number of years ${n}_{\rm year}$ over which a sufficient number $\overline{n}_{\rm large}$ of large events are gathered, as can be seen by combining (12) and (37):
${\rm RSE}_{\rm SAL}=\frac{\sqrt{2}}{\sqrt{\overline{n}_{\rm large}}}=\frac{\sqrt{2}}{\sqrt{f_{\rm large}^{\rm all}n_{\rm year}}}$ (38)
Hence, to ensure for SALEDI a statistical accuracy ${\rm RSE}_{\rm SAL}^{\rm max}$ so that ${\rm RSE}_{\rm SAL}\leq{\rm RSE}_{\rm SAL}^{\rm max}$, the minimum number of integer years required is
$n_{\rm year}=\text{round up to next integer}\Big[\frac{2}{f_{\rm large}^{\rm all}({\rm RSE}_{\rm SAL}^{\rm max})^{2}}\Big]$ (39)
In this paper, we adopt ${\rm RSE}_{\rm SAL}^{\rm max}\approx 0.1$, implying that the standard deviation of an estimate of SALEDI is no more than 10% of its mean value. Under reasonable normality assumptions, this further indicates that $\sim$68% of the estimates of SALEDI lie within 10% of the mean, and that $\sim$90% of the estimates of SALEDI lie within 16% of the mean. The $n_{\rm year}$ and $\overline{n}_{\rm large}$} values for the five utilities based on ${\rm RSE}_{\rm SAL}^{\rm max}\approx 0.1$ are given in Table I. The value of $n_{\rm year}$ implies tracking SALEDI with an $n_{\rm year}$ year sliding window. This is practical for the five utilities in Table I that have $n_{\rm year}=$ 2, 3, or 5.
@@ VI Conclusions
Quantifying distribution system resilience in terms of the observed customer minutes interrupted per customer served (CMIp) in large events can be considered one of the basic challenges for resilience metrics. However, detailed outage data from five utilities in the USA shows heavy-tailed distributions and high variability in event CMIp. Indeed these observed data show evidence of very heavy tails (slope magnitude $\alpha<1.5$ of an exceedance function on a log-log plot.) A consequence of the heavy tails is that simply adding up or averaging the CMIp for large events does not give practical indices. These impractical indices (such as SPALEDI, SPALED, and conventional CVAR) vary erratically and their relative standard errors are generally so large that too many (one or two orders of magnitude more) observed large events would be needed to enable usable statistical accuracy.
@@
SALEDI is the product of the annual frequency of large events $f_{\rm large}$ and the ALED metric that is an average magnitude of large events. More precisely, ALED is the mean of the logarithms of normalized large event CMIp. Risk is often described by exceedance functions, and resilience risk describes the tails of these functions. We show that ALED estimates the slope magnitude $\alpha$ of the tail of the probability exceedance function of CMIp on a log-log plot. Thus, ALED quantifies the trend of the observed large blackouts, which is significant because this trend may well govern future blackouts of high risk that are larger than those already observed. Our definition of large events is chosen so that the large events lie in the approximately linear tail region of the log-log plot to ensure that ALED, and in turn SALEDI, describes this trend. This approach requires a suitable approximately linear tail region for the largest observed blackouts, as shown in the data for the five utilities.
The paper also systematically interprets and compares the metrics SALEDI and ALED, which can be considered to be areas under the tails of exceedance functions. ALED is a version of CVAR in which the value is the logarithm of normalized CMIp. SAIDI is the area under a frequency exceedance function for outage CMIp.
In summary, since the CMIps in large events have high variability and heavy-tailed distributions, simply adding up or taking the mean of CMIp in large events does not produce practical indices. A logarithmic transformation of CMIp converts heavy tails to light tails for which the usual statistics and indices work. Thus the overall conclusion is simple: Based on five cases of distribution utility data in the USA, CMIp in large events should be measured on a logarithmic scale, so that practical resilience indices tracking customer minutes interrupted in large events should be calculated using the logarithm of the CMIp.
@@
The truncated lognormal fit is quadratic on a log-log plot, but with a large enough $\sigma$ so that the quadratic coefficient $1/(2\sigma^{2})$ is small [16, section 1.2.3] and the linear approximation is acceptable. The standard deviation parameter $\sigma$ of the best-fit truncated lognormal distribution for the utilities 1,2,3,4,5 are 4.05, 4.53, 3.80, 4.16, 1.39, respectively.
@@
Let $N_{\rm large}$ be Poisson[$\overline{n}_{\rm large}$], assumed independent of i.i.d. samples $X_{1}$, $X_{2}$, …, $X_{N_{\rm large}}$ of random variable $X$. Let
@@
${\rm E}N_{\rm large}={\rm Var}N_{\rm large}=\overline{n}_{\rm large}$. A Wald equation gives ${\rm E}S={\rm E}[N_{\rm large}{\rm E}X]/n_{\rm year}$ and a Blackwell-Girshick equation gives
$~\hskip-14.22636pt{\rm\hskip-14.22636ptVar}S=\,$ $({\rm E}[N_{\rm large}]{\rm Var}X+{\rm Var}[N_{\rm large}]({\rm E}X)^{2}])/n_{\rm year}^{2}$
$=\,$ $\overline{n}_{\rm large}({\rm Var}X+({\rm E}X)^{2})/n_{\rm year}^{2}\text{ and}$
${\rm RSE}[S]$ $=\frac{\sigma[S]}{{\rm E}S}=\frac{\sqrt{1+{\rm Var}X/({\rm E}X)^{2}}}{\sqrt{\overline{n}_{\rm large}}}=\frac{\sqrt{1+{\rm RSE}[X]^{2}}}{\sqrt{\overline{n}_{\rm large}}}$
Also ${\rm E}S_{\rm mean}={\rm E}X$ and ${\rm Var}S_{\rm mean}=({\rm Var}X)/\overline{n}_{\rm large}$ and
${\rm RSE}[S_{\rm mean}]$ $=\frac{\sigma[S_{\rm mean}]}{{\rm E}S_{\rm mean}}=\frac{\sigma[X]}{{\rm E}X\sqrt{\overline{n}_{\rm large}}}=\frac{{\rm RSE}[X]}{\sqrt{\overline{n}_{\rm large}}}.$

### v2

@@ IV-A Why the logarithm in SALEDI is needed
is a Seemingly Plausible Average Large Event Duration Index that is the mean number of normalized CMIp for large events. SPALED is ALED without the logarithm. We now quantify the variability of SPLEDI and SPALED to show that practical estimates of SPLEDI and SPALED require too many samples.
@@
Let $Y$ be the random variable for the normalized event CMIp $M/M_{\rm large}$. We will take $Y\sim P_{b}$ or $Y\sim LN_{b}$. According to (23) and (21), SPALED is the mean of $N_{\rm large}$ samples of $Y$, where $N_{\rm large}$ has mean $n_{\rm large}$, and SPLEDI is SPALED multiplied by annual frequency $f_{\rm large}$. Then, using Appendix B, the relative standard error of SPLEDI is
${\rm RSE}_{\rm SPL}$ $=\frac{\sqrt{1+({\rm RSE}_{Y})^{2}}}{\sqrt{n_{\rm large}}}$ (26)
Then, similarly to section V-B, where we take ${\rm RSE}_{\rm SAL}=0.1$, the minimum number of large events to have ${\rm RSE}_{\rm SPL}=0.1$ is
$n_{{\rm large}\,Y}^{\rm minSPL}=100(1+({\rm RSE}_{Y})^{2})$ (27)
For approximating the bounded distribution of CMIp with $Pb$, Table I shows that utilities 1,2,3,4 have $n_{{\rm large}\,Pb}^{\rm minSPL}$ at least two orders of magnitude greater than $n_{\rm large}^{\rm min}=200$. That is, for the same statistical accuracy, SPLEDI requires at least two orders of magnitude more large events than SALEDI, making it impractical as a resilience metric calculated from observed data. For example, if SALEDI requires 2 years of data to accumulate enough large events for reasonable statistical accuracy, then SPLEDI requires at least 200 years. Note that the ratio $n_{{\rm large}\,Pb}^{\rm minSPL}/n_{\rm large}^{\rm min}$ is $1+({\rm RSE}_{Pb})^{2}]/2$, which is independent of the choice of statistical accuracy ${\rm RSE}_{\rm SAL}=0.1$. For utility 5, $n_{{\rm large}\,Pb}^{\rm minSPL}=18\,000$ which is almost two order of magnitudes greater than $n_{\rm large}^{\rm min}=200$ and so requires more large events by a factor of $90$.
For approximating the bounded distribution of CMIp with $LNb$, Table I shows that utilities 1,2,3 have $n_{{\rm large}LNb}^{\rm minSPL}$ two orders of magnitude greater than $n_{\rm large}^{\rm min}=200$, so that SPLEDI requires at least two orders of magnitude more large events than SALEDI for the same accuracy. Utility 4 requires more large events by a factor of $38$, and utility 5 requires more large events by a factor of $4$.
Similarly, using Appendix B, the relative standard error of SPALED is ${\rm RSE}_{\rm SPA}={\rm RSE}_{Y}/\sqrt{n_{\rm large}}$ so that the minimum number of large events to have ${\rm RSE}_{\rm SPA}=0.1$ is $n_{{\rm large}\,Y}^{\rm minSPA}=100({\rm RSE}_{Y})^{2}$, and the ratio $n_{{\rm large}\,Y}^{\rm minSPA}/n_{\rm large}^{\rm min}=({\rm RSE}_{Y})^{2}$. Then $n_{{\rm large}\,Y}^{\rm minSPA}$ for both $Y=Pb$ and $Y=LNb$ is two orders of magnitude larger than ALED’s $n_{\rm large}^{\rm minALE}=100$ (obtained starting from (13)) except that $n_{{\rm large}LNb}^{\rm minSPA}$ = 7600 and 700 for utilities 4 and 5 respectively. Even if it could in theory be computed from large quantities of data, SPALED, since it is a mean, cannot be interpreted as a typical value due to the heavy tails.
@@
IV-B Comparing the uses of SALEDI and SAIDI
SAIDI without major event days and SALEDI quantify different aspects of CMIp and are complementary. SAIDI without major event days focuses on the CMIp of common individual outages and quantifies reliability by excluding major event days, whereas SALEDI focuses on the CMIp of large events and quantifies resilience.
SAIDI with major event days measures the total impact of all sustained outages experienced by customers in a year. Although SAIDI with major event days is some sort of combination of reliability and resilience, it varies erratically as it is mostly driven by the very large variability of the large events that can occur in each year. Moreover, it is desirable to separately monitor reliability and resilience with different metrics. While improvements to reliability can often also improve resilience or not affect resilience, they are distinct because the scale and urgency and risk of large events differ so much from routine outages. Therefore proposed upgrades should be evaluated separately for reliability and resilience.
@@ IV-C CVAR interpretations
Therefore ALED, which works well as an index, is a logarithmic form of CVAR [^11] However, ${\rm CVAR}_{\rm lnM}$ differs from ALED in the choice of the threshold $M_{\rm large}$ for large events. For ${\rm CVAR}_{\rm lnM}$, $M_{\rm large}$ is VAR, the value at a given quantile of normalized event CMIp, whereas for ALED as presented here, $M_{\rm large}$ is the lowest point of the linear region of the exceedance function.. Moreover, (5) shows that
@@ IV-D Heavy tails and extrapolating the trend of extreme events
The bounded Pareto distribution (24) linearly extrapolates the slope of the exceedance curve tail beyond the maximum observed normalized CMIp $M_{\rm maxobs}/M_{\rm large}$ up to the maximum normalized CMIp $M_{\rm max}/M_{\rm large}$ for the largest possible blackout. This allows the resilience to be evaluated while including the unobserved future blackouts of the highest customer impact. The linear extrapolation of the trend is somewhat uncertain, but the extrapolation nevertheless seems more reasonable than ignoring the even more extreme events indicated by the observed trend—it seems unlikely that the largest possible blackout has already been observed in the data. It is the case that sums of samples of heavy-tailed distributions are dominated by the maximum sample that occurs [^12] This is called a “catastrophe principle” and proved in [15, chapter 3].. For example, both SAIDI with major event days and SPLEDI add up normalized CMIp, and suffer high variability due to this effect. High-impact blackouts with extremely high CMIp values are rare but inevitable due to heavy tails. This emphasizes the importance, despite its inherent uncertainties, of trying to estimate the extrapolated trend governing the most extreme events to encompass these extreme events in the resilience index. The need to obtain this linear extrapolation is one reason that we confine the large events to the linear region of the tail. We assume the existence of a suitable linear region across a sufficient range of the largest observed blackouts because it is present in the data we analyze here.
@@ V Practical details of SALEDI
V-A Choosing the large event threshold $M_{\rm large}$
SALEDI is defined over events whose CMIp $M_{i}$ exceeds the $M_{\rm large}$ threshold. It is intended that $M_{\rm large}$ is calculated once and then stays fixed thereafter. We want to select an $M_{\rm large}$ that satisfies the following criteria:
  • $M_{\rm large}$ should be small enough to give sufficient large events to keep the variability ${\rm RSE}_{\rm SAL}$ sufficiently small.
@@ V-A Choosing the large event threshold $M_{\rm large}$
    Select all values above $M_{0}$, i.e., $S_{0}=\{M_{i}:M_{i}\geq M_{0}\}$
@@
    $\alpha_{0}=(1/|S_{0}|)\sum_{i=1}^{|S_{0}|}{\rm ln}\{M_{i}/M_{0}:M_{i}\geq M_{0}\}$
    where $|S_{0}|$ is the number of values in the selected data.
@@
    Calculate Kolmogorov-Smirnov distance $D$ between the empirical CDF of the selected data $F_{S_{0}}(M)$ and the CDF $F_{P_{0}}(M)$ of a Pareto distribution with parameter $\alpha_{0}$: $D=\underset{M_{i}\geq M_{0}}{\rm max}|F_{S_{0}}(M_{i})-F_{P_{0}}(M_{i})|$
@@
V-B Number of large events $n_{\rm large}^{\rm min}$ required for accuracy
The statistical variability of SALEDI is governed by the number of large events $n_{\rm large}$ used, as can be seen from (12). Hence, rearranging (12), to ensure for SALEDI a statistical accuracy ${\rm RSE}_{\rm SAL}^{\rm max}$ so that ${\rm RSE}_{\rm SAL}\leq{\rm RSE}_{\rm SAL}^{\rm max}$, the minimum number of large events required is
$n_{\rm large}^{\rm min}$ $=2/({\rm RSE}_{\rm SAL}^{\rm max})^{2}$ (35)
Suppose that $n_{\rm large}^{\rm all}$ large events are observed in all the outage data over $n_{\rm year}^{\rm all}$ years. Then the average annual frequency of large events using all the observed years of data is
$f_{\rm large}^{\rm all}=n_{\rm large}^{\rm all}/n_{\rm year}^{\rm all}$ (36)
and the minimum number of years of data required to ensure for SALEDI a statistical accuracy ${\rm RSE}_{\rm SAL}^{\rm max}$ is
$n_{\rm year}^{\rm min}=n_{\rm large}^{\rm min}/f_{\rm large}^{\rm all}$ (37)
In this paper, we adopt ${\rm RSE}_{\rm SAL}^{\rm max}=0.1$, implying that the standard deviation of an estimate of SALEDI is no more than 10% of its mean value. Then (35) gives $n_{\rm large}^{\rm min}=200$.
Under reasonable normality assumptions, ${\rm RSE}_{\rm SAL}^{\rm max}=0.1$ further indicates that $\sim$68% of the estimates of SALEDI lie within 10% of the mean, and that $\sim$90% of the estimates of SALEDI lie within 16% of the mean.
The $n_{\rm year}$ values for the five utilities based on ${\rm RSE}_{\rm SAL}^{\rm max}=0.1$ are given in Table I. The value of $n_{\rm year}^{\rm min}$ implies tracking SALEDI with an $n_{\rm year}^{\rm min}$ year sliding window. This is practical for the five utilities in Table I that have $n_{\rm year}^{\rm min}\approx$ 2, 3, or 5.
@@ VI Conclusions
Quantifying distribution system resilience in terms of the observed customer minutes interrupted per customer served (CMIp) in large events can be considered one of the basic challenges for resilience metrics. However, detailed outage data from 5 utilities in the USA shows heavy-tailed distributions and high variability in event CMIp. Indeed these observed data show evidence of very heavy tails (slope magnitude $\alpha<1.5$ of an exceedance function on a log-log plot). A consequence of the heavy tails is that simply adding up or averaging the CMIp for large events does not give practical indices. These impractical indices (such as SPALEDI, SPALED, and conventional CVAR) vary erratically and their relative standard errors are usually so large that too many (one or two orders of magnitude more) observed large events would be needed to enable usable statistical accuracy.
@@
SALEDI is the product of the annual frequency of large events $f_{\rm large}$ and the ALED metric that is an average magnitude of large events. More precisely, ALED is the mean of the logarithms of normalized large event CMIp. Risk is often described by exceedance functions, and resilience risk describes the tails of these functions. ALED-1 estimates the slope magnitude $\alpha$ of the tail of the probability exceedance function of CMIp on a log-log plot. Thus, ALED quantifies the trend of the observed large blackouts, which is significant because this trend may well govern future blackouts of high risk that are larger than those already observed. Our definition of large events is chosen so that the large events lie in the approximately linear tail region of the log-log plot to ensure that ALED, and in turn SALEDI, describes this trend. This approach requires a suitable approximately linear tail region for the largest blackouts, as observed in the data for the 5 utilities. SALEDI is normalized by the number of customers served, so it is comparable across utilities of different sizes.
The paper also systematically interprets and compares the metrics SALEDI and ALED, which can be considered to be areas under the tails of exceedance functions. ALED is a version of CVAR in which the value is the logarithm of normalized event CMIp. SAIDI is the area under a frequency exceedance function for outage CMIp.
Resilience mitigation should decrease the frequency and severity of customer impacts, and in this regard, $f_{\rm large}$, ALED, and SALEDI usefully measure different aspects. $f_{\rm large}$ is the frequency of large events, which is driven by the extremes of weather exceeding a given customer impact. ALED measures an average (on a log scale) large event impact for the customers, which results from a combination of weather severity, infrastructure hardness, and restoration speed. A utility with high $f_{\rm large}$ but low ALED is being hit frequently but withstands and recovers well relative to the weather severities. SALEDI combines ALED and $f_{\rm large}$ to measure the total annual large event customer impact. SALEDI gives a single value for optimization and tracking resilience over the years, that can play for resilience for customers the same role as SAIDI does for reliability. A utility would first compute the large event threshold $M_{\rm large}$ and the required $n_{\rm year}^{\rm min}$ once from its full historical dataset. It would then maintain a rolling $n_{\rm year}^{\rm min}$ window and update SALEDI annually to track resilience.
In summary, since the CMIps in large events have high variability and heavy-tailed distributions, simply adding up or taking the mean of CMIp in large events does not produce practical indices. A logarithmic transformation of CMIp converts heavy tails to light tails for which the usual statistics and indices work. Thus an overall conclusion is simple: Based on five cases of distribution utility data in the USA, CMIp in large events should be measured on a logarithmic scale, so that practical resilience indices tracking customer minutes interrupted in large events should be calculated using the logarithm of the CMIp.
@@
The truncated lognormal fit is quadratic on a log-log plot, but with a large enough $\sigma$ so that the quadratic coefficient $1/(2\sigma^{2})$ is small [15, section 1.2.3] and the linear approximation is acceptable. The standard deviation parameter $\sigma$ of the best-fit truncated lognormal distribution for the utilities 1,2,3,4,5 are 4.05, 4.53, 3.80, 4.16, 1.39, respectively.
@@
Let $N_{\rm large}$ be Poisson[$n_{\rm large}$], assumed independent of i.i.d. samples $X_{1}$, $X_{2}$, …, $X_{N_{\rm large}}$ of random variable $X$. Let
@@
${\rm E}N_{\rm large}={\rm Var}N_{\rm large}=n_{\rm large}$. A Wald equation gives ${\rm E}S=n_{\rm large}{\rm E}X/n_{\rm year}$ and a Blackwell-Girshick equation gives
$~\hskip-14.22636pt{\rm\hskip-14.22636ptVar}S=\,$ $({\rm E}[N_{\rm large}]{\rm Var}X+{\rm Var}[N_{\rm large}]({\rm E}X)^{2})/n_{\rm year}^{2}$
$=\,$ $n_{\rm large}({\rm Var}X+({\rm E}X)^{2})/n_{\rm year}^{2}\text{ and}$
${\rm RSE}_{S}$ $=\frac{\sigma[S]}{{\rm E}S}=\frac{\sqrt{1+{\rm Var}X/({\rm E}X)^{2}}}{\sqrt{n_{\rm large}}}=\frac{\sqrt{1+({\rm RSE}_{X})^{2}}}{\sqrt{n_{\rm large}}}$
Also ${\rm E}S_{\rm mean}={\rm E}X$ and ${\rm Var}S_{\rm mean}=({\rm Var}X)/n_{\rm large}$ and
${\rm RSE}_{S_{\rm mean}}$ $=\frac{\sigma[S_{\rm mean}]}{{\rm E}S_{\rm mean}}=\frac{\sigma[X]}{{\rm E}X\sqrt{n_{\rm large}}}=\frac{{\rm RSE}_{X}}{\sqrt{n_{\rm large}}}.$